OpenStudy (kweb2017):

Given the slope of the line and one point that it passes through, represent the equation of the line in point-slope form. Then rewrite the equation in slope-intercept form. (-4,7); m= -3/4

pooja195 (pooja195):

y=mx+b slope is known y=-3/4x+b plug in the coordinates 7=-3/4(-4)+b solve for b

OpenStudy (kweb2017):

how...

pooja195 (pooja195):

Cancel out the 4's -4+4=0 Now you have 7=3+b subtract 3 from both sides b=?

OpenStudy (kweb2017):

4

pooja195 (pooja195):

Good that should give you your equation.